a train moving at a constant speed travels at 16.5 kilometers in 5 minutes what is the distance in 1 (sec)

Answer by srinivas.g(540)

train covers 16.5 km in 5 minutes
it means it covers 16.5 km in 300 seconds ( 5 min = 5*60 =300 seconds)
In one second it can cover = 16.5 Km over 300
= 0.055 km
= 0.055 *1000 meters
= 55 m
distance in 1 sec = 55 m